An object is launched at 45ᵒ to the horizontal on the level ground. What is the range of the projectile if it has an initial velocity of 180 ft/s? Neglect air resistance.


1
Expert's answer
2021-05-18T13:00:09-0400

We can find the range of the projectile from the formula:


"R=\dfrac{v^2sin2\theta}{g},""R=\dfrac{(180\ \dfrac{ft}{s})^2sin2\cdot45^{\circ}}{32.17\ \dfrac{ft}{s^2}}=1007\ ft."